OpenStudy (volleyballlover55):

Compare 8 ⋅ 10^7 to 2 ⋅ 10^4. 8 ⋅ 10^7 is 4,000 times larger than 2 ⋅ 10^4. 8 ⋅ 10^7 is 400 times larger than 2 ⋅ 10^4. 8 ⋅ 10^7 is 40 times larger than 2 ⋅ 10^4. 8 ⋅ 10^7 is 4 times larger than 2 ⋅ 10^4.

OpenStudy (anthonyym):

Well first you can compare the 8 and 2, how many times 8 is greater than 2. Then you compare the 10 to the power part and see how many times 10^7 is greater than 10^4 (do 10^7 divided by 10^4).

OpenStudy (anonymous):

what is 80,000000/20,000?

OpenStudy (anthonyym):

Because you're really doing \[\frac{ 8*10^7 }{ 2*10^4 }=\frac{ 8 }{ 2 }*\frac{ 10^7 }{ 10^4 }\]

OpenStudy (anthonyym):

So knowing that what do you get?

OpenStudy (volleyballlover55):

@surjithayer

OpenStudy (volleyballlover55):

b?

OpenStudy (anonymous):

\[\frac{ 8*10^7 }{ 2*10^4 }=4*10^7*10^{-4}=4*10^{7-4}=4*10^3=4*1000=4000\]

OpenStudy (volleyballlover55):

oh.. so A?

OpenStudy (anonymous):

correct.

OpenStudy (volleyballlover55):

ok next :) Convert 3,200,000 to scientific notation